# Structure-Aligned Protein Language Model

### v1

@@ Abstract
Protein language models (pLMs) pre-trained on vast protein sequence databases excel at various downstream tasks but lack the structural knowledge essential for many biological applications. To address this, we integrate structural insights from pre-trained protein graph neural networks (pGNNs) into pLMs through a latent-level contrastive learning task. This task aligns residue representations from pLMs with those from pGNNs across multiple proteins, enriching pLMs with inter-protein structural knowledge. Additionally, we incorporate a physical-level task that infuses intra-protein structural knowledge by optimizing pLMs to predict structural tokens. The proposed dual-task framework effectively incorporates both inter-protein and intra-protein structural knowledge into pLMs. Given the variability in the quality of protein structures in PDB, we further introduce a residue loss selection module, which uses a small model trained on high-quality structures to select reliable yet challenging residue losses for the pLM to learn. Applying our structure alignment method to the state-of-the-art ESM2 and AMPLIFY results in notable performance gains across a wide range of tasks, including a $12.7\%$ increase in ESM2 contact prediction. The data, code, and resulting SaESM2 and SaAMPLIFY models will be released on Hugging Face.
@@ 1 Introduction
Building on recent progress in natural language processing (Brown et al., 2020; Devlin et al., 2019), researchers have focused on pre-training protein language models (pLMs) on vast databases of protein sequences with masked language modeling (Rives et al., 2019; Hayes et al., 2024; Fournier et al., 2024) and next token prediction (Ferruz et al., 2022). These pLMs produce representations that demonstrate substantial potential across a variety of biological applications, including protein function annotation, enzyme-catalyzed reaction prediction, and protein classification (Hu et al., 2022).
While the sequence-only nature of pLMs contributes to their widespread adoption, they often struggle in tasks requiring detailed structural insights. For instance, ESM-GearNet outperforms ESM2 by $9.7\%$ on the Human Protein-Protein Interaction classification task (Xu et al., 2022; Su et al., 2024). In this paper, we aim to develop a pLM that preserves its sequence-only nature for broader applicability yet is augmented with structural insights.
Given the availability of open-source pre-trained protein graph neural networks (pGNNs) (Zhang et al., 2023; Chen et al., 2023; Jumper et al., 2021), we investigate integrating pGNN-derived structural insights into pLMs. Specifically, we introduce a latent-level contrastive learning task for the structural alignment of pLMs. As illustrated in Figure 1, this task aligns residue hidden representations from the pLM ($\bm{h}_{a}$) with those from the pGNN ($\bm{h}_{g}$) across a batch of $B$ proteins. During this process, the pGNN is frozen while the pLM is optimized to minimize the contrastive learning loss, enriching the pLM with inter-protein [^1] Note that “inter-protein” and “intra-protein” refer to tasks involving multiple proteins and within a single protein, respectively. This usage differs from the biological definition, where “inter-protein” refers to interactions between two proteins, and “intra-protein” refers to interactions within a single protein chain. structural knowledge. However, pure contrastive alignment may overemphasize residue-level patterns across the dataset, neglecting intra-protein11footnotemark: 1 structural context (Zheng and Li, 2024). To address this, we introduce a physical-level task that trains the pLM to predict structural tokens $\bm{z}$ (representing physical conformations (van Kempen et al., 2022)) from its residue representations $\bm{h}_{a}$. This reinforces the encoding of each residue within its protein, thereby enriching the pLM with intra-protein structural knowledge.
@@
We conducted $10$ ablations to validate our design choices. Our analysis demonstrates that the proposed dual-task framework is crucial for optimal performance, with residue loss selection providing further gains. The models were evaluated on $6$ tasks from xTrimoPGLM (Chen et al., 2024), $9$ from SaProt (Su et al., 2023), and pseudo-perplexity using the high-quality held-out validation set from Fournier et al. (2024).
@@
  • We propose a dual-task framework that integrates inter-protein and intra-protein residue-level structural knowledge into pLMs.
@@
  • We conduct extensive experiments that demonstrate the effectiveness of our method.
@@ 2 Preliminaries
In this section, we introduce the preliminaries of protein language models, structure embeddings, and structure tokens used in this study, and provide a more detailed review of related work in Appendix B.
@@ 2.1 Protein Language Models
Proteins can be represented as sequences of amino acids, where each amino acid $a_{i}$ belongs to the set of the $20$ common types. A protein sequence of length $L$ is denoted as $\bm{a}=(a_{1},a_{2},\ldots,a_{L})$.
@@
[…]
@@ 2.2 Protein Structure Embeddings
Given a protein graph $\bm{g}$, where each residue is a node and edges are defined based on both sequential and structural distances, and a pre-trained protein GNN model, the model outputs residue-level embeddings $\text{pGNN}(\bm{g})\in\mathbb{R}^{L\times D_{g}}$, where $L$ is the number of residues, and $D_{g}$ is the embedding dimension of the graph-based residue representation.
@@ 2.3 Protein Structure Tokens
Foldseek (van Kempen et al., 2022) introduces an efficient method for tokenizing protein structures, where each residue $i$ is described by its geometric conformation relative to its spatially closest residue $j$. While this approach has significantly accelerated homology detection, it incurs substantial information loss, thereby limiting its applicability for tasks requiring detailed structural reconstruction. To address this limitation, ProToken (Lin et al., 2023a) employs a symmetric encoder-decoder architecture that enables high-fidelity reconstruction of protein structures from tokens. Despite this advancement, these tokens have shown limited effectiveness in broader downstream applications Zhang et al. (2024a).
Recently, Hayes et al. (2024) developed an effective vector quantization variational autoencoder (VQ-VAE) tokenizer and integrated structure and sequence into a multi-modal large language model called ESM3. This approach effectively combines both modalities, improving the model’s versatility. We did not compare to ESM3 due to licensing restrictions. AIDO (Zhang et al., 2024a) further enhances structure tokenization by introducing a novel VQ-VAE with an equivariant encoder and an invariant decoder, ensuring a more robust representation of protein structures.
@@ 3.1 Dual-Task Framework
where $s$ follows the approach in CLIP (Radford et al., 2021). In our experiments, we primarily use GearNet (Zhang et al., 2023) as the pGNN, pre-trained on the AlphaFold2 database (Varadi et al., 2022). We also evaluate the Evoformer within AlphaFold2 (Jumper et al., 2021) but find GearNet to be more effective for our purpose.
@@ 3.2 Residue Loss Selection
To address the challenge posed by ambiguous or inaccurate protein structures in the PDB (Burley et al., 2019), we propose a residue loss selection module. This module ensures both effectiveness and efficiency prioritizes residue losses that align with high-quality protein structures.
@@
We begin by curating a high-quality reference set using resolution and R-free metrics (Morris et al., 1992). Structures with resolution below $2.0$Å and R-free lower than $0.20$ are selected as a clean reference set. We then train a smaller language model, either ESM2 150M or AMPLIFY 120M, on the reference set with the same loss in Equation 9 and denote the optimized reference model parameters as $\bm{\theta}^{r}$, $\bm{\alpha}^{r}$, $\bm{W}^{r}$, $\bm{\beta}^{r}$. The resulting reference model is used to assess the residue loss of the alignment corpus.
@@
For each residue loss discussed in subsection 3.1, we compute the excess loss, defined as the difference between the residue loss of the current model and that of the reference model:
@@
where $\mathbbm{1}\big{(}\mathcal{L}_{\text{a2g}}(\Delta,i,b_{1}),\rho\big{)}$ equals $1$ if $\mathcal{L}_{\text{a2g}}(\Delta,i,b_{1})$ ranks in the top $\rho$ of all $\mathcal{L}_{\text{a2g}}(\Delta,i,b_{1})$ values, and 00 otherwise. This selection process is applied similarly for the other two types of losses. By focusing on challenging yet reliable residue losses, the residue loss selection module improves overall training effectiveness and efficiency.
@@ 4.1 Structure Alignment Details
Our alignment dataset is sourced from the PDB database, comprising $129,732$ proteins, including $116,713$ proteins designated for training and $13,019$ for validation/reference. The training protocol is adapted from the AMPLIFY stage-2 configuration (Fournier et al., 2024) with several modifications. We extend the pre-training with $20$ epochs on our alignment dataset, with the learning rate linearly warming up from 00 to the peak rate over the first two epochs, followed by a cosine decay schedule for the subsequent $18$ epochs. The peak learning rate for the language model is set at $1\times 10^{-4}$, as per the AMPLIFY standard, while other modules, such as the structural linear classifier and the contrastive learning module, are set at $1\times 10^{-3}$. The selection ratio $\rho$ is set as $0.8$. We employ the Zero Redundancy Optimizer (ZeRO) with DeepSpeed for efficiency and use $8$ H100 GPUs. The effective batch size is $4,096$ samples at a sequence length of $2,048$, with longer proteins being randomly truncated.
@@ 4.2 Baseline Models
We evaluate the following sequence-only baseline pLMs: (1) ESM2: the standard ESM2 650M model (Lin et al., 2022); (2) AMPLIFY: the standard AMPLIFY 350M model (Fournier et al., 2024); (3) ESM2-s: a variant of ESM2 fine-tuned for fold classification (Zhang et al., 2024b); (4) ISM: a variant of ESM2 optimized for structure token prediction (Ouyang-Zhang et al., 2024). We denote our structure-aligned ESM2 and AMPLIFY models as SaESM2 and SaAMPLIFY.
4.3 Structure Prediction
@@
Structure-aligned models are expected to capture more nuanced insights of protein structures. We consider the following structure prediction tasks from xTrimoPGLM Chen et al. (2024): (1) Contact: two residues are considered in contact if their $C_{\beta}$ atoms lie within 8Å (Rao et al., 2019). We evaluate this task using Top L/5 accuracy, considering residue pairs with a sequence separation greater than $6$ and a sequence length cutoff of $512$. (2) Fold Classification (Fold): classify each protein sequence into one of $1,195$ fold classes (Hou et al., 2018), with accuracy as the evaluation metric. (3) Secondary Structure (SS): assign each residue to one of three secondary structure types (Rao et al., 2019), using accuracy as the evaluation metric.
To assess the quality of the learned representations, we freeze the backbone model and train a linear head for $20$ epochs using a batch size of $128$. We use a learning rate of $1\times 10^{-3}$, with betas set to $(0.9,0.95)$ and a weight decay of $0.01$ (Fournier et al., 2024). The linear head has a hidden size of $128$, following (Chen et al., 2024). The linear head operates on residue embeddings for the token-level task (SS), on the mean-pooled residue embedding for the sequence-level task (Fold), and on pairwise residue embedding for the Contact task. We further visualize residue embeddings with secondary structure labels to assess structural alignment effectiveness in Appendix C.
@@
As shown in Table 2, SaESM2 and SaAMPLIFY outperform their respective base models on all structure prediction tasks as well as existing alignment baselines on two out of three tasks, improving Contact P@L/5 by $13\%$ for ESM2 and $42\%$ for AMPLIFY. ESM2-s, directly trained on fold classification, outperforms our alignment method on the corresponding task.
Table 1: Results on structure prediction.
Contact Fold SS
P@L/5 ($\uparrow$) Acc ($\uparrow$) Acc ($\uparrow$)
ESM2 $54.14$ $56.13$ $81.19$
ESM2-s $57.38$ 67.26 $79.87$
ISM $55.55$ $47.87$ $83.24$
SaESM2 (ours) 61.02 $60.73$ 84.91
AMPLIFY $26.31$ $33.63$ $79.86$
SaAMPLIFY (ours) $37.34$ $49.01$ $84.12$
Table 2: Results on mutation effect prediction.
Fluorescence Fitness (GB1) Stability
Sp. ($\uparrow$) Sp. ($\uparrow$) Sp. ($\uparrow$)
ESM2 $0.687$ $0.916$ $0.742$
ESM2-s $0.688$ $0.942$ $0.726$
ISM $0.694$ $0.933$ $0.759$
SaESM2 0.695 0.951 0.787
AMPLIFY $0.684$ $0.922$ $0.734$
SaAMPLIFY $0.694$ $0.924$ $0.756$
4.4 Mutation Effect Prediction

### v2

@@ Abstract
Protein language models (pLMs) pre-trained on vast protein sequence databases excel at various downstream tasks but often lack the structural knowledge essential for some biological applications. To address this, we introduce a method to enrich pLMs with structural knowledge by leveraging pre-trained protein graph neural networks (pGNNs). First, a latent-level contrastive learning task aligns residue representations from pLMs with those from pGNNs across multiple proteins, injecting inter-protein structural information. Additionally, a physical-level task integrates intra-protein information by training pLMs to predict structure tokens. Together, the proposed dual-task framework effectively incorporates both inter- and intra-protein structural knowledge into pLMs. Given the variability in the quality of protein structures in PDB, we further introduce a residue loss selection module that uses a small model trained on high-quality structures to select reliable yet challenging residue losses for the pLM to learn. Applying our structure alignment method as a simple, lightweight post-training step to the state-of-the-art ESM2 and AMPLIFY yields notable performance gains. These improvements are consistent across a wide range of tasks, including substantial gains in deep mutational scanning (DMS) fitness prediction and a 59% increase in P@L for ESM2 650M contact prediction on CASP16. Furthermore, we demonstrate that these performance gains are robust, scaling with model sizes from 8M to 650M and extending to different downstream tasks.
@@ 1 Introduction
Building on recent progress in natural language processing (Brown et al., 2020; Devlin et al., 2019), researchers have focused on pre-training protein language models (pLMs) on vast databases of protein sequences with masked language modeling (Rives et al., 2019; Hayes et al., 2024; Fournier et al., 2024) and next token prediction (Ferruz et al., 2022). These pLMs learn representations that researchers have demonstrated hold substantial potential across a variety of biological applications, including protein function annotation, enzyme-catalyzed reaction prediction, and protein classification (Hu et al., 2022). Additionally, Rives et al. (2019) observed that structural information emerged in the models’ latent representations without supervision. Nonetheless, while the sequence-only nature of pLMs contributes to their widespread adoption, they often struggle in tasks requiring detailed structural insights. For instance, the structured informed ESM-GearNet outperforms ESM2 by $9.7\%$ on the Human Protein-Protein Interaction classification task (Xu et al., 2022; Su et al., 2024). In this paper, we aim to develop a pLM that preserves its sequence-only nature for broader applicability yet is augmented with structural insights.
Given the availability of open-source pre-trained protein graph neural networks (pGNNs) (Zhang et al., 2023; Chen et al., 2023; Jumper et al., 2021), we investigate integrating pGNN-derived structural insights into pLMs. Specifically, we introduce a latent-level contrastive learning task for the structural alignment of pLMs. As illustrated in Figure 1, this task aligns residue hidden representations from the pLM ($\bm{h}_{a}$) with those from the pGNN ($\bm{h}_{g}$) across a batch of $B$ proteins. During this process, the pGNN is frozen while the pLM is optimized to minimize the contrastive learning loss, enriching the pLM with inter-protein [^1] Note that “inter-protein” and “intra-protein” refer to tasks involving multiple proteins and within a single protein, respectively. This usage differs from the biological definition, where “inter-protein” refers to interactions between two proteins, and “intra-protein” refers to interactions within a single protein chain. structural knowledge. However, pure contrastive alignment may overemphasize differing residue-level patterns across the dataset, neglecting intra-protein11footnotemark: 1 structural context (Zheng and Li, 2024). To address this, we add a physical-level task that trains the pLM to predict structural tokens $\bm{z}$ (representing physical conformations (van Kempen et al., 2022)) from its residue representations $\bm{h}_{a}$. This reinforces the encoding of each residue within its protein, thereby enriching the pLM with intra-protein structural knowledge.
@@
We conducted $10$ ablations to validate our design choices. Our analysis demonstrates that the proposed dual-task framework improves performance, with residue loss selection providing further gains. The models were evaluated on a comprehensive suite of benchmarks. We assess performance on deep mutational scanning (DMS) fitness prediction using ProteinGym (Notin et al., 2023) and on direct structural validation using a contact prediction task on withheld data from CASP16 (Yuan et al., 2025). We further test generalization on $9$ tasks from xTrimoPGLM (Chen et al., 2024) and $9$ from SaProt (Su et al., 2023), and evaluate language modeling fidelity using pseudo-perplexity on the high-quality held-out validation set from Fournier et al. (2024). We find that structure alignment is a computationally lightweight post-training step that, depending on the downstream task, either matches or exceeds the performance of the original model.
@@
  • We propose a dual-task framework that integrates inter-protein and intra-protein residue-level structural knowledge into pLMs, while fully retaining their language modeling capabilities.
@@
  • We conduct extensive experiments that demonstrate the effectiveness of our method across model sizes, model families and downstream tasks.
@@ 2 Preliminaries
In this section, we introduce the preliminaries of protein language models, structure embeddings, and structure tokens used in this study, and provide a more detailed review of related work in Appendix A.
@@ 2.1 Protein Language Models
Proteins can be represented as sequences of amino acids, where each amino acid $a_{i}$ belongs to the set of $20$ common types. A protein sequence of length $L$ is denoted as $\bm{a}=(a_{1},a_{2},\ldots,a_{L})$.
@@
A pre-trained pLM parameterized by $\bm{\theta}$ is represented as $\text{pLM}(\cdot;\bm{\theta})$. The latent representation of a protein sequence $\bm{a}$ is denoted as $\text{pLM}(\bm{a};\bm{\theta})\in\mathbb{R}^{L\times D_{a}}$, where $D_{a}$ is the embedding dimension. During pre-training, a subset of positions $\mathcal{M}\subset\{1,\ldots,L\}$ is replaced with a $\text{mask}]$ token:
@@ 2.2 Protein Structure Embeddings
Given a protein graph $\bm{g}$, where each residue is a node and edges are defined based on both sequential and structural distances, a pre-trained protein GNN model outputs residue-level embeddings $\text{pGNN}(\bm{g})\in\mathbb{R}^{L\times D_{g}}$, where $L$ is the number of residues, and $D_{g}$ is the embedding dimension of the graph-based residue representation.
@@ 2.3 Protein Structure Tokens
Foldseek (van Kempen et al., 2022) introduces an efficient method for tokenizing protein structures, where each residue $i$ is described by its geometric conformation relative to its spatially closest residue $j$. While this approach has significantly accelerated homology detection, it incurs substantial information loss, thereby limiting its applicability to tasks requiring detailed structural reconstruction. To address this limitation, ProToken (Lin et al., 2023a) employs a symmetric encoder-decoder architecture that enables high-fidelity reconstruction of protein structures from tokens. Despite this advancement, these tokens have shown limited effectiveness in broader downstream applications (Zhang et al., 2024a).
Recently, Hayes et al. (2024) developed an effective vector quantization variational autoencoder (VQ-VAE) tokenizer and integrated structure and sequence into a multi-modal protein language model called ESM3. This approach effectively combines both modalities, improving the model’s versatility. While we could not evaluate ESM3 ourselves due to licensing restrictions, we were able to retrieve its reported performance on the ProteinGym benchmark. AIDO (Zhang et al., 2024a) further enhances structure tokenization by introducing a novel VQ-VAE with an equivariant encoder and an invariant decoder, ensuring a more robust representation of protein structures.
@@ 3.1 Dual-Task Framework
where $s$ follows the approach in CLIP (Radford et al., 2021). The contrastive step is similar to that of Robinson et al. (2023), with three notable exceptions: we keep the pGNN frozen, because our objective is to align the pLM, they discard the language modeling head, which is contrary to our final loss Equation 5 and they discard our physical level task (subsection 3.1), replacing it with another, higher-level, inter-protein contrastive loss. Their objective also seems to be more the improvement of protein sequential models, built in complete isolation from a folding task.
In our experiments, we primarily use GearNet (Zhang et al., 2023) as the pGNN, pre-trained on the AlphaFold2 database (Varadi et al., 2022). We also evaluated the Evoformer within AlphaFold2 (Jumper et al., 2021) but found GearNet embeddings to be more effective for our purpose.
@@
The proposed method is, to the best of our knowledge, the first to incorporate MLM regularization. Consequently, among similarly structure-aligned or contrastively fine-tuned sequence-only pLMs, we are the first to be evaluated on language modeling downstream tasks, such as Deep Mutation Scanning, which are crucial in drug discovery pipelines. Also, contrary to existing models such as SaProt (Su et al., 2024), there is no need for structure as input to enrich residue embeddings.
This structure-agnostic capability is essential, given that proteins with intrinsically disordered regions, which lack a fixed tertiary structure, constitute a significant portion of the proteome. Independence from structural input ensures the model’s applicability to any protein, including those with uncharacterized structures or those for which in silico folding predictions may be unreliable.
@@ 3.2 Residue Loss Selection
To address the challenge posed by ambiguous or inaccurate protein structures in the PDB (Burley et al., 2019), we propose a residue loss selection module. This module ensures both effectiveness and efficiency by prioritizing residue losses that align with high-quality protein structures.
@@
We begin by curating a high-quality reference set using resolution and R-free metrics (Morris et al., 1992). Structures with resolution below 2.02.0Å and R-free lower than $0.20$ are selected as a clean reference set. We then train a smaller language model on the reference set with the same loss in Equation 9 and denote the optimized reference model parameters as $\bm{\theta}^{r}$, $\bm{\alpha}^{r}$, $\bm{W}^{r}$, $\bm{\beta}^{r}$. The resulting reference model is used to assess the residue loss of the alignment corpus.
@@
For each residue loss discussed in §3.1, we compute the excess loss, defined as the difference between the residue loss of the current model and that of the reference model:
@@
where $\mathbbm{1}\big(\mathcal{L}_{\text{a2g}}(\Delta,i,b_{1}),\rho\big)$ equals $1$ if $\mathcal{L}_{\text{a2g}}(\Delta,i,b_{1})$ ranks in the top $\rho$ of all $\mathcal{L}_{\text{a2g}}(\Delta,i,b_{1})$ values, and 0 otherwise. This selection process is applied similarly for the other two types of losses. By focusing on challenging yet reliable residue losses, the residue loss selection module improves overall training effectiveness and efficiency.
@@ 4.1 Structure Alignment Details
We aligned ESM2 and AMPLIFY using $129,732$ proteins from OpenFold (Ahdritz et al., 2023) present in the PDB database, of which $116,713$ are for training and $13,019$ for validation. We systematically verified that our sequences were deposited in the PDB in December 2021 to the latest. As a consequence, our training set is fully deduplicated against CASP16 , meaning CASP16 represents a gold standard test set for downstream evaluation of our models.
The training protocol is adapted from the AMPLIFY stage-2 configuration (Fournier et al., 2024) with several modifications. We extend the pre-training with $20$ epochs on our alignment dataset, with the learning rate linearly warming up from 0 to the peak rate over the first two epochs, followed by a cosine decay schedule for the subsequent $18$ epochs. The peak rate for the language model is set at $1\times 10^{-4}$, as per the AMPLIFY standard, while other modules, such as the structural linear classifier and the contrastive learning module, are set at $1\times 10^{-3}$. The selection ratio $\rho$ is set to $0.8$.
We employ the Zero Redundancy Optimizer (ZeRO) with DeepSpeed and use $8$ H100 GPUs. The effective batch size is $4,096$ samples at a sequence length of $2,048$, with longer proteins being randomly truncated. Our post-training alignment method is particularly compute efficient, taking under 6 hours for the largest ESM2 model considered, and under 1 hour for the smallest model.
@@ 4.2 Baseline Models
We evaluate the following sequence-only baseline pLMs: (1) ESM2: the standard ESM2 650M model (Lin et al., 2022); (2) AMPLIFY: the standard AMPLIFY 350M model (Fournier et al., 2024); (3) ESM2-S: a variant of ESM2 fine-tuned for fold classification (Zhang et al., 2024b); (4) ISM: a variant of ESM2 optimized for structure token prediction (Ouyang-Zhang et al., 2024); (5) S-PLM: a different contrastive post-training method applied to ESM2 (Wang et al., 2025) [^2] Note that S-PLM has approximately 100M additional parameters compared to ESM2 and all the other baselines.. We denote our structure-aligned ESM2 and AMPLIFY models as SaESM2 and SaAMPLIFY.
4.3 Supervised Downstream Task Performance
To evaluate the effectiveness of our structure alignment, we benchmark our models against their un-aligned counterparts on a comprehensive suite of supervised downstream tasks. For these tasks, the pLM is fine-tuned, either with a head and frozen backbone or with full-model fine-tuning, for each specific objective. We group these into structural property prediction, supervised mutation effect prediction, and broader functional property prediction.
Whenever possible we report either confidence intervals at 95% computed with bootstrapping for downstream property prediction tasks, or the full distributions over the 217 sub-benchmark assays of ProteinGym (see §4.4).
4.3.1 Structural Property Prediction
@@
To test the hypothesis that structure-aligned models capture more nuanced insights of protein structures, we evaluate on the following structure prediction tasks from xTrimoPGLM (Chen et al., 2024): (1) Contact: two residues are considered in contact if their $C_{\alpha}$ atoms lie within 8Å (Rao et al., 2019). We evaluate this task using Top L/5 precision, as Rives et al. (2021), considering residue pairs with a sequence separation greater than $6$ and a sequence length cutoff of $512$. In order to compare existing models without data leakage, we select the subset of CASP16 proteins that have already been deposited in PDB and contain at least one long-range contact. We also report the original xTrimoPGLM test split (created from the trRosetta dataset), and (2) Fold Classification (Fold): classify each protein sequence into one of $1,195$ fold classes (Hou et al., 2018), with accuracy as the evaluation metric. (3) Secondary Structure (SS): assign each residue to one of three secondary structure types (Rao et al., 2019), using accuracy as the evaluation metric.
To assess the quality of the learned representations, we freeze the backbone model and train a linear head for $20$ epochs using a batch size of $128$. We use a learning rate of $1\times 10^{-3}$, with betas set to $(0.9,0.95)$ and a weight decay of $0.01$ (Fournier et al., 2024). The linear head has a hidden size of $128$, following the methodology of xTrimoPGLM. The linear head operates on residue embeddings for the token-level task (SS), on the mean-pooled residue embedding for the sequence-level task (Fold), and on pairwise residue embedding for the Contact task. We further visualize residue embeddings with secondary structure labels to assess structural alignment effectiveness in Appendix B.
@@
As shown in Table 1, SaESM2 and SaAMPLIFY outperform their respective base models on all structure prediction tasks as well as existing alignment baselines on two out of three tasks, improving Contact P@L/5 on CASP16 by $59\%$ for ESM2 and $15\%$ for AMPLIFY. This is a direct validation of the way inter- and intra-protein structural knowledge is infused by our method. Due to the fact that ESM2-S was directly trained on fold classification with unfrozen backbone, it outperforms our alignment method on the corresponding task.
In Appendix C.1, we show that these conclusions still hold across model size and families for secondary structure prediction and, to a lesser extent, fold classification, providing proof that the method scales.
Table 1: Results on supervised downstream tasks. We report the primary metric for each task. Values are formatted as Metric [95% Confidence Interval]. The best-performing model within each family (ESM2-based and AMPLIFY-based) is in bold. Models within the best Confidence Interval are in italic.
Contact (P@L/5 $\uparrow$) Fold $S$ Fitness Stability
Model trRosetta CASP16 Acc ($\uparrow$) Acc ($\uparrow$) Sp. ($\uparrow$) Sp. ($\uparrow$)
ESM2 0.390 [0.379 8120.379\,812, 0.400 2240.400\,224] 0.181 [0.145 5440.145\,544, 0.223 5430.223\,543] 0.677 [0.661 5290.661\,529, 0.691 7390.691\,739] 0.845 [0.843 4430.843\,443, 0.846 7220.846\,722] 0.945 [0.937 0280.937\,028, 0.951 4070.951\,407] 0.744 [0.735 8750.735\,875, 0.752 1260.752\,126]
ESM2-S 0.387 [0.377 1620.377\,162, 0.397 6260.397\,626] 0.182 [0.147 8170.147\,817, 0.221 6680.221\,668] 0.764 [0.749 6920.749\,692, 0.778 0590.778\,059] 0.811 [0.809 1570.809\,157, 0.812 6600.812\,660] 0.961 [0.955 4910.955\,491, 0.965 4530.965\,453] 0.765 [0.756 0440.756\,044, 0.772 6470.772\,647]
ISM 0.426 [0.416 5050.416\,505, 0.435 9510.435\,951] 0.220 [0.180 7150.180\,715, 0.261 6540.261\,654] 0.598 [0.580 4560.580\,456, 0.614 0570.614\,057] 0.840 [0.838 4510.838\,451, 0.841 9150.841\,915] 0.957 [0.950 6260.950\,626, 0.961 7720.961\,772] 0.558 [0.544 5330.544\,533, 0.571 2650.571\,265]
S-PLM 0.403 [0.393 5070.393\,507, 0.413 4340.413\,434] 0.229 [0.190 3070.190\,307, 0.272 7440.272\,744] 0.662 [0.646 4240.646\,424, 0.676 6410.676\,641] 0.821 [0.819 1190.819\,119, 0.822 7020.822\,702] 0.947 [0.939 7620.939\,762, 0.952 4010.952\,401] 0.661 [0.650 8670.650\,867, 0.671 8970.671\,897]
SaESM2 0.461 [0.450 4930.450\,493, 0.471 2270.471\,227] 0.288 [0.250 3750.250\,375, 0.327 0210.327\,021] 0.681 [0.664 9200.664\,920, 0.696 3630.696\,363] 0.865 [0.863 0090.863\,009, 0.866 0280.866\,028] 0.957 [0.951 4930.951\,493, 0.961 7630.961\,763] 0.820 [0.812 6190.812\,619, 0.827 3480.827\,348]
AMPLIFY 0.253 [0.244 6780.244\,678, 0.261 7500.261\,750] 0.155 [0.120 3880.120\,388, 0.191 6460.191\,646] 0.487 [0.470 3990.470\,399, 0.502 1660.502\,166] 0.811 [0.809 3960.809\,396, 0.812 9350.812\,935] 0.947 [0.941 2420.941\,242, 0.952 5260.952\,526] 0.713 [0.704 1260.704\,126, 0.722 0850.722\,085]
SaAMPLIFY 0.320 [0.310 5600.310\,560, 0.328 1390.328\,139] 0.169 [0.143 9820.143\,982, 0.194 6930.194\,693] 0.576 [0.557 3370.557\,337, 0.592 7870.592\,787] 0.849 [0.846 9500.846\,950, 0.850 2450.850\,245] 0.948 [0.941 3120.941\,312, 0.953 4420.953\,442] 0.747 [0.738 5960.738\,596, 0.755 7130.755\,713]
Figure 2: Longe range contact prediction precision. (Left) Results on the CASP16 test set. (Right) Results on the trRosetta test split from the xTrimoPGLM evaluation pipeline. Structure-aligned models significantly outperform their baseline on both test sets, starting at the 150M parameters model size. Note that the confidence intervals on the harder split (CASP16) are larger due to a small sample size.
4.3.2 Functional Property Prediction
We evaluate SaESM2 and SaAMPLIFY on a broad suite of downstream property prediction tasks (Xu et al., 2022; Dallago et al., 2021), which rely on structural information to some extent but are not direct structure prediction tasks. These include predictions of thermostability, metal ion binding, protein localization (DeepLoc), enzyme commission numbers (EC), gene ontology annotations (GO), and protein–protein interactions (HumanPPI) for tasks and evaluation pipeline from (Su et al., 2024).
In addition to downstream structural evaluations (§4.3.1) and supervised mutation effect prediction (§4.3.3) from xTrimoPGLM, we further evaluate on 3 other properties: enzyme catalytic efficiency, peptide-MHC/TCR binding affinity and peptide-HLA-MHC affinity. In all cases, we follow the data splits and training protocols from the respective papers.
As detailed in §C.1 and §C.2, we observed no meaningful change in performance between the structure-aligned models and their respective unaligned baselines, leading us to conclude that the proposed method does not degrade functional property prediction. Notably, the confidence intervals of the 3 structure‑aligned models (SaESM2, ESM2‑S, and ISM) fully overlap with those of their baseline across all 9 SaProt tasks. Moreover, out of the 27 points of comparison, the structure‑aligned models outperform their baselines only 14 times, barely over half, suggesting that structure alignment does not provide meaningful prediction improvement or degradation on these tasks. We offer in §C.2 three possible hypotheses for this.
4.3.3 Supervised Mutation Effect Prediction